### v1

# Mechanical Squeezed-Fock Qubit: Towards Quantum Weak-Force Sensing

Abstract
Mechanical qubits offer unique advantages over other qubit platforms, primarily in terms of coherence time and possibilities for enhanced sensing applications, but their potential is constrained by the inherently weak nonlinearities and small anharmonicity of nanomechanical resonators. We propose to overcome this shortcoming by using squeezed Fock states of phonons in a parametrically driven nonlinear mechanical oscillator. We find that, under two-phonon driving, squeezed Fock states become eigenstates of a Kerr-nonlinear mechanical oscillator, featuring an energy spectrum with exponentially enhanced and tunable anharmonicity, such that the transitions to higher energy states are exponentially suppressed. This enables us to encode the mechanical qubit within the ground and first excited squeezed Fock states of the driven mechanical oscillator. This kind of mechanical qubit is termed mechanical squeezed-Fock qubit. We also show that our mechanical qubit can serve as a quantum sensor for weak forces, with its resulting sensitivity increased by at least one order of magnitude over that of traditional mechanical qubits. The proposed mechanical squeezed-Fock qubit provides a powerful quantum phonon platform for quantum sensing and information processing.
Introduction.—Mechanical resonators in the quantum regime represent a pivotal platform for quantum science and technologies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], offering promising pathways toward scalable quantum computation and ultra-precise sensing [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Compared to their electromagnetic counterparts, mechanical resonators offer longer lifetimes, greater compactness [22], and the capability for direct coupling to different degrees of freedom [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Such devices enable quantum squeezing of mechanical motion [39, 40, 41, 42, 43], quantum transduction [44, 45, 46, 47], and quantum entanglement [48, 49, 50, 51, 52, 53]. In particular, mechanical resonators exhibiting nonlinearities have been viewed as intriguing candidates for the realization of mechanical qubits in quantum computation [54, 55, 56]. Furthermore, their ability to couple to a variety of external weak forces, including gravitational forces [57, 58, 59], magnetic forces, or others, enables quantum force sensing with substantially improved sensitivity [60, 61, 62, 63].
Traditionally, a mechanical qubit is encoded in the first two Fock states ($|0\rangle$ and $|1\rangle$) of a Kerr-nonlinear oscillator [54, 55, 56]. Coherent control via external fields requires the energy difference $\alpha$ between the $|0\rangle\rightarrow|1\rangle$ and $|1\rangle\rightarrow|2\rangle$ transitions to greatly exceed the phonon decoherence rate ($\alpha\gg\gamma$). This condition ensures quantum operations excite state $|2\rangle$ negligibly. However, the development in mechanical qubits has been fundamentally limited by the inherent physical constraints of nanomechanical resonators—notably weak nonlinearities and insufficient anharmonicity—which hinder robust state engineering and coherent control. This limitation could be overcome by coupling mechanical resonators to other strongly nonlinear quantum systems such as superconducting qubits [64, 65, 66, 67, 68, 69, 70, 71]. However, this may induce additional decoherence channels inherited from the auxiliary nonlinear quantum systems [54, 55, 56], which in turn limit the performance of mechanical qubits for quantum sensing and computation. Thus, a natural question arises: Can we find a new mechanism to boost mechanical oscillators with weak Kerr nonlinearities into high-fidelity qubits for quantum sensing with unprecedented sensitivity?
In this work, we provide an answer to the above question. We find that squeezed Fock states are the eigenstates of a Kerr-nonlinear mechanical oscillator under two-phonon driving, with a spectrum exhibiting exponentially enhanced and tunable anharmonicities. We show that the transitions to higher energy states are exponentially suppressed, and the dynamics of the system is confined to the subspace spanned by the first and second squeezed Fock states. In this way, a mechanical qubit can be encoded in the first two squeezed-Fock phonon states of the drivien Kerr-nonlinear oscillator, which is termed mechanical squeezed-Fock qubit. Compared to traditional mechanical Fock qubits, this squeezed-Fock qubit proposal offers critical advantages: (i) Under weak Kerr nonlinearity (small $K$), it exponentially suppresses leakage to higher states by enhancing anharmonicity from $\alpha\sim K$ to $\alpha\sim Ke^{4r}$ (where $r$ is the two-phonon driving parameter), overcoming the inherent limitation of conventional designs. (ii) It maintains qubit validity across arbitrary regimes of Kerr nonlinearity—including $K>\gamma$, $K\sim\gamma$, and $K<\gamma$. This enables direct transformation of weakly anharmonic mechanical oscillators into robust qubits, dramatically simplifying experimental implementation by eliminating the need for strong intrinsic nonlinearity.
We employ this mechanical squeezed-Fock qubit for weak force sensing. In stark contrast to traditional mechanical Fock-qubit protocols [62, 55], the sensitivity for weak force detection increases by one or two orders of magnitude. Our mechanical squeezed-Fock qubit provides a powerful new platform for quantum sensing and computation.
Figure 1: Energy Level diagram and exponentially enhanced anharmonicity. (a) Schematic of a doubly clamped mechanical resonator with weak Kerr nolinear. Applying a two-phonon drive, the system behaves as high-fidelity qubits rather than resonators. (b) Spectrum of a mechanical qubit without and with a two-phonon drive in the condition of $K\sim\gamma_{0}$ (not to scale). The nonlinear terms $K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}$ causes the harmonic potential of the mechanical resonator to become anharmonic, but the weak $K$ is not sufficient to select two distinct states as a qubit. The two-phonon drive makes the nonlinear intensity be enhanced exponentially with a scale of $\sim e^{4r}$, such that the two lowest state (red dashed square) can be chosen as a qubit. (c) The exponentially enhanced values of $\alpha$, $\gamma$ and their ratio are shown on a logarithmic scale as functions of the parameter $r$.
Mechanical squeezed-Fock qubit.—Conventionally, a mechanical resonator [as shown in Fig. 1(a)], whose eigenstates are the phonon Fock states $|n\rangle$ with eigenenergy $E_{n}$, is characterized by an equally spaced energy levels. Introducing a Kerr nonlinearity breaks this harmonicity, resulting in unevenly spaced energy levels. Since the Kerr nonlinear term commutes with the number operator, the eigenstates remain Fock states $|n\rangle$. Consequently, the lowest two states can be selected to form an effective qubit. The Hamiltonian for this system is ($\hbar=1$)
$\hat{H}=\omega_{a}\hat{a}^{\dagger}\hat{a}+K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a},$ (1)
where $\omega_{a}$ is the resonator frequency, $\hat{a}^{\dagger}$ ($\hat{a}$) is the phonon creation (annihilation) operator, and $K$ quantifies the strength of the phonon-phonon interaction. The dynamics of the system is therefore governed by the Lindblad master equation:
$\frac{d\rho}{dt}=-i[\hat{H},\rho]+\frac{\gamma_{0}}{2}\mathcal{D}_{\hat{a},\hat{a}^{\dagger}}\rho,$ (2)
where $\rho$ is the system’s density operator, $\gamma_{0}$ is the decoherence rate, and $\mathcal{D}_{\hat{A},\hat{B}}\rho\equiv 2\hat{A}\rho\hat{B}-\hat{B}\hat{A}\rho-\rho\hat{B}\hat{A}$. Here, the anharmonicity $\alpha$ is introduced to assess whether the resonator can function as a mechanical qubit. This anharmonicity is defined as the energy difference between the $|1\rangle\rightarrow|2\rangle$ and $|0\rangle\rightarrow|1\rangle$ transitions: $\alpha_{0}=2K$. For a mechanical qubit, the anharmonicity must significantly exceed the decoherence rate ($\alpha\gg\gamma_{0}$). This ensures that the system’s energy predominantly oscillates between the qubit states $|0\rangle$ and $|1\rangle$ with minimal leakage to $|2\rangle$ under the coherent quantum operations. However, achieving strong nonlinearity (large $K$) without auxiliary nonlinear quantum systems experimentally remains challenging. For weak $K$, the small detuning between the $\{|0\rangle,|1\rangle\}$ and $\{|1\rangle,|2\rangle\}$ transitions leads to significant leakage to higher states, which compromises the validity of the two-level approximation. While leakage can be suppressed by using a sufficiently weak drive, this approach renders the qubit highly vulnerable to decoherence, thereby preventing its use in quantum tasks.
Figure 2: Validity of mechanical squeezed Fock qubit. (a) and (b) are the population $P_{n}$ of Fock states $|n\rangle$ in the absence and presence of decoherence. (c) and (d) are the population $P_{n}$ of squeezed Fock states $|n\rangle_{S}$ in the absence and presence of decoherence. (e) and (f) are the Wigner functions of phonon modes at three points during a direct Fock qubit and squeezed Fock qubit Rabi oscillation, respectively. Here, the initial parameters $\gamma_{0}=K$ in (b) and (d), $r=1.5$ in (c), (d) and (f).
We propose to achieve an exponentially enhanced anharmonicity by applying a two-phonon drive to the resonator. Specifically, we consider pumping the system with a periodic drive of frequency $\omega_{p}/2$ and amplitude $\Omega_{p}$. In the rotating frame defined by the unitary transformation $\hat{U}_{0}(t)=\exp(-i\omega_{p}\hat{a}^{\dagger}\hat{a}t)$, the Hamiltonian (1) becomes
$\hat{H}=\delta_{a}\hat{a}^{\dagger}\hat{a}+K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\frac{\Omega_{p}}{2}\left[e^{-i\theta}\hat{a}^{2}+e^{i\theta}(\hat{a}^{\dagger})^{2}\right],$ (3)
where $\delta_{a}=\omega_{a}-\omega_{p}$ is the detuning and $\theta$ is the drive phase. The Hamiltonian (3) can be diagonalized via the Bogoliubov transformation $\hat{b}=\cosh r\cdot\hat{a}+e^{i\theta}\sinh r\cdot\hat{a}^{\dagger}=\hat{S}\hat{a}\hat{S}^{\dagger}$, with the squeezing operator $\hat{S}=\exp\left(r^{*}\hat{a}^{2}/2-r\hat{a}^{\dagger 2}/2\right)$ and the squeezing parameter $r$ satisfies $\tanh(2r)=\Omega_{p}/\delta_{a}$. The diagonalized Hamiltonian is given by
$\hat{H}_{\text{eff}}=\omega_{b}\hat{b}^{\dagger}\hat{b}+U_{b}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b},$ (4)
where $\omega_{b}=\sqrt{\delta_{a}^{2}-\Omega_{p}^{2}}+K(8\cosh^{2}r\sinh^{2}r+4\sinh^{4}r)$ represents the squeezed-resonator frequency and $U_{b}=[3\cosh(4r)+1]K/4$ is the enhanced nonlinear strength. Obviously, the Bogoliubov transformation yields a new mechanical mode $\hat{b}$ with frequency $\omega_{b}$ and exponentially strengthened nonlinear interaction $U_{b}\sim e^{4r}K$. In addition, the two-phonon drive transforms the system’s eigenstates from Fock states $|n\rangle$ to squeezed Fock states $|n\rangle_{S}\equiv\hat{S}|n\rangle$, as illustrated in Fig. 1(b). Selecting the two lowest squeezed states $|0\rangle_{S}$ and $|1\rangle_{S}$ as qubit basis states yields the anharmonicity
$\alpha=2U_{b}=\frac{3\cosh(4r)+1}{2}K,$ (5)
which demonstrates an exponential enhancement proportional to $e^{4r}$. The system dynamics is therefore governed by the master equation:
$\frac{d\rho}{dt}=$ $-i[\hat{H}_{\text{eff}},\rho]+\frac{\gamma_{0}}{2}(\mathcal{N}+1)\mathcal{D}_{\hat{b},\hat{b}^{\dagger}}\rho+\frac{\gamma_{0}}{2}\mathcal{N}\mathcal{D}_{\hat{b}^{\dagger},\hat{b}}\rho$ (6)
$-\frac{\gamma_{0}}{2}\mathcal{M}\mathcal{D}_{\hat{b},\hat{b}}\rho-\frac{\gamma_{0}}{2}\mathcal{M}^{*}\mathcal{D}_{\hat{b}^{\dagger},\hat{b}^{\dagger}}\rho,$
where $\mathcal{N}=\sinh^{2}r$ and $\mathcal{M}=e^{-i\theta}\cosh r\sinh r$. The first term governs coherent dynamics, while the second and third terms represent incoherent dissipation: $\mathcal{D}_{\hat{b},\hat{b}^{\dagger}}\rho$ and $\mathcal{D}_{\hat{b}^{\dagger},\hat{b}}\rho$ correspond to energy emission and absorption processes, respectively. The final two terms describe two-phonon correlated processes induced by the two-phonon drive [72, 73, 74, 75]. Here $\mathcal{N}$ and $\mathcal{M}$ are squeezing parameters with amplitude $r$ and phase angle $\theta$, where $\theta=0$ and $\theta=\pi$ correspond to the squeezing of position and momentum space, respectively.
Note that the two-phonon drive not only exponentially enhances the anharmonicity but also amplifies noise. However, the anharmonicity enhancement scales as $\sim e^{4r}$, significantly exceeding the noise amplification at a rate $\sim e^{2r}$. This contrast is illustrated in Fig. 1(c). Initially assuming the decoherence rate $\gamma_{0}$ exceeds the bare anharmonicity $\alpha_{0}$ by an order of magnitude ($\gamma_{0}\sim 10\alpha_{0}$), we observe both quantities grow exponentially with increasing $r$. Crucially, since the growth rate of $\alpha$ is $e^{2r}$ times that of $\gamma\equiv\gamma_{0}\sinh^{2}r$, $\alpha$ surpasses $\gamma$ near $r=1$ and rapidly dominates at larger $r$. We further quantify this advantage through the phonon anharmonicity-decoherence ratio $\alpha/\gamma$ versus $r$. These results demonstrate that weak intrinsic nonlinearities ($\alpha_{0}\ll\gamma_{0}$) can be exponentially enhanced via two-phonon driving. This enables robust encoding of a mechanical qubit in the two lowest squeezed Fock states $|0\rangle_{S}$ and $|1\rangle_{S}$.
To intuitively demonstrate the validity of the mechanical qubit encoded in the squeezed Fock states $|0\rangle_{S}$ and $|1\rangle_{S}$, Fig. 2 compares the the populations as well as the Wigner function of phonons without (Fock states) and with (squeezed Fock state) the two-phonon drive, respectively. Under coherent driving at rate $\Omega_{d}$ [Fig. 2(a) and (b) in the absence and presence of decoherence], the populations of Fock states $|2\rangle$ (blue triangles) and $|3\rangle$ (orange inverted triangles) exhibit significant temporal evolution. In contrast, Fig. 2(c) demonstrates complete suppression of leakage to the squeezed Fock states $|2\rangle_{S}$ and $|3\rangle_{S}$, whose populations remain negligible. This confirms that the two-phonon drive confines the dynamics of the system to the qubit subspace, enabling predominant oscillations between $|0\rangle_{S}$ and $|1\rangle_{S}$. When we take the decoherence into consideration, as shown in Fig. 2(d), we can obtain a damped oscillation between the two squeezed-Fock states $|0\rangle_{S}$ and $|1\rangle_{S}$. From the perspective of the Wigner function, we can perform the qubit operations and visualize the resulting states, as shown in Fig. 2(e) and 2(f). We assume that the initial state is $|0\rangle$ and $|0\rangle_{S}$, corresponding to the cases without and with two-phonon drive, respectively. For the case of the former, we can find that the mechanical system gradually evolves to the state with a sub-Poissonian distribution. By contrast, the state under two-phonon drive evolves approximately to $|1\rangle_{S}$ at $\Omega_{d}t=\pi$, and eventually returns to $|0\rangle_{S}$. This further confirms that two-phonon driving enables weakly nonlinear mechanical resonators to behave as high-fidelity qubits rather than resonators.
Figure 3: The validity of mechanical squeezed Fock qubit in the Ramsey sensing. (a) Flowchart of the Ramsey sensing. (b) The population of $|1\rangle_{S}$ obtained by the Ramsey sensing and (c) The sensitivity $\delta k_{\text{min}}/k_{0}$ versus nonlinear strength $K$ and parameter $r$. Here, $\Omega_{\pi/2}$ is the driving strength of the $\pi/2$ pulse and the initial parameter $\gamma_{0}=\Omega_{\pi/2}$.
Sensing of the spring constant of restoring forces.—We have thus established a mechanical qubit encoded in the squeezed Fock states $|0\rangle_{S}$ and $|1\rangle_{S}$. We now consider utilizing this mechanical squeezed Fock qubit to sense the weak force field, with the coupling term given by $H_{V}=kx^{2}/2$, and $k$ the effective spring constant induced by the external signal. Specifically, the qubit Hamiltonian is given by $\hat{H}_{0}=\omega_{b}\hat{\sigma}_{z}/2$, where ${\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z}}$ denote the Pauli operators. The projection of signal Hamiltonian into the Hilbert space spanned by $\{|0\rangle_{S},\,|1\rangle_{S}\}$ can be expressed as $\hat{H}_{V}=\frac{1}{2}kx_{0}^{2}e^{2r}\hat{\sigma}_{z}$. Here, $x_{0}=1/\sqrt{2m\omega_{a}}$ represents the zero-point fluctuation amplitude of the resonator with mass $m$. The total Hamiltonian is then
$\hat{H}_{\text{tot}}^{\text{V}}=\frac{1}{2}(\omega_{b}+\omega_{V})\hat{\sigma}_{z},$ (7)
where $\omega_{V}\equiv kx_{0}^{2}e^{2r}$. The signal induces a frequency shift in the qubit, which can be directly measured via Ramsey interferometry [key steps are shown in Fig. 3(a)] [62]. Under dissipative decoherence, the encoding process is governed by the master equation:
$\frac{d\rho}{dt}=$ $-i[\hat{H}_{\text{tot}}^{\text{V}},\rho]+\frac{\gamma_{0}}{2}(\mathcal{N}+1)\mathcal{D}_{\hat{\sigma}_{-},\hat{\sigma}_{+}}\rho+\frac{\gamma_{0}}{2}\mathcal{N}\mathcal{D}_{\hat{\sigma}_{+},\hat{\sigma}_{-}}\rho$ (8)
$+\gamma_{0}\mathcal{M}\hat{\sigma}_{-}\rho\hat{\sigma}_{-}-\gamma_{0}\mathcal{M}^{*}\hat{\sigma}_{+}\rho\hat{\sigma}_{+},$
where $\hat{\sigma}_{-}=\hat{\sigma}_{+}^{\dagger}=|0\rangle_{S}\langle 1|$ are the ladder operators of qubit. This equation describes the dynamics of two-level system under a squeezed thermal reservoir and we have used the condition: $\hat{\sigma}_{-}\hat{\sigma}_{-}=\hat{\sigma}_{+}\hat{\sigma}_{+}=0$ [76]. We assume the initial state is $|0\rangle_{S}$. After the initial $\pi/2$ pulse, free evolution and a second $\pi/2$ pulse, the system evolves into [77]
$\rho(t)$ $=\frac{1}{2}[(1+\text{Re}[c(t)])\,|0\rangle_{S}\langle 0|+(1-\text{Re}[c(t)])\,|1\rangle_{S}\langle 1|]$ (9)
[…]
where $c(t)=\exp\{[-\gamma_{0}\cosh(2r)/2-i(\omega_{b}+\omega_{V})]t\}$, $\kappa=1/\cosh 2r$. The probability $P_{|1\rangle_{S}}$ of measuring the qubit in the excited state is then given by
$P_{|1\rangle_{S}}=\frac{1}{2}-\frac{1}{2}e^{-\gamma_{0}\cosh(2r)t/2}\cos[(\omega_{b}+\omega_{V})t],$ (10)
[…]
We simulate the $P_{|1\rangle_{S}}$ during the Ramsey measurement process under different parameters in Fig. 3(b). In the absence of two-phonon drive, i.e., $r=0$, weak Kerr nonlinearity $K$ causes significant leakage to higher states, leading to $P_{|1\rangle_{S}}<0.5$. Here, the observed oscillations arise from the interference between primary and leakage evolution paths [76]. With the introduction of the two-phonon drive, we can find that both energy leakage and parasitic oscillations are exponentially suppressed. In addition, the quantity of interest here is the difference $\delta P_{|1\rangle_{S}}$ between the probability measured with and without the signal and we can express it as [76]
$\delta P_{|1\rangle_{S}}=\pm\frac{1}{2}e^{-\gamma_{0}\cosh(2r)t/2}x_{0}^{2}e^{2r}\delta kt,$ (11)
where $\delta k$ denotes the spring constant variation. We can find the difference $\delta P_{|1\rangle_{S}}$ grows linearly with evolution time $t$. It indicates that longer $t$ means more sensitivity to changes in the external field (higher resolution), but it is also subject to the qubit decoherence.
We next estimate the fundamental sensitivity through signal-to-noise ratio (SNR) analysis, which is defined as [62]
$\text{SNR}=\frac{\delta P_{|1\rangle_{S}}}{\sigma_{p}}=e^{-\gamma_{0}\cosh(2r)t/2}x_{0}^{2}e^{2r}\delta kC\sqrt{N}t$ (12)
where $\sigma_{p}=1/2C\sqrt{N}$ is the total noise with the number of experiment $N=T/(t+t_{m})$, $T$ is the total available measurement time and $t_{m}$ is the extra time needed to initialize, manipulate, and read out the sensor, $C$ represents the efficiency parameter of readout. We take the value $C=1$ and $t_{m}=0$. The sensitivity is then defined as the minimum detectable signal $\delta k_{\text{min}}$ that yields unit SNR for an integration time of $T=1$[55]. Its minimal over the encoding time is
$\delta k_{\text{min}}=\frac{\sqrt{\gamma_{0}\cosh(2r)e}}{x_{0}^{2}e^{2r}}\approx\frac{\sqrt{\gamma_{0}e}}{x_{0}^{2}e^{r}},$ (13)
which is obtained in the condition of $d\delta k_{\text{min}}/dt=0$, i.e., $t=1/[\gamma_{0}\cosh(2r)]$. The sensitivity is exponentially promoted by a factor $\sim e^{r}$ compared to conventional mechanical Fock qubits. We have shown the sensitivity $\delta k_{\text{min}}/k_{0}$ in Fig. 3(c), where $k_{0}=\sqrt{\gamma_{0}e}/x_{0}^{2}$ is the standard sensitivity of $kx^{2}/2$ sensing by the mechanical Fock qubits [76]. Here, The red dashed line indicates the reference $\delta k_{\text{min}}/k_{0}=1$. We can clearly see that $\delta k_{\text{min}}/k_{0}$ decreases as $r$ increases, which indicates that our proposal provides higher precision for quantum sensing.
Experimental feasibility.—To explore the experimental feasibility of our proposal, we here discuss the relevant parameters. We consider a mechanical resonator with frequency $\omega_{a}/2\pi=600$ MHz and decoherence $\gamma_{0}/2\pi=3$ kHz, which results in a quality factor $Q=2\times 10^{5}$. For a finite temperature of 10 mK, the average thermal phonon numbers $\overline{n}\approx 0.00595\ll 1$ and can be ignored. Assuming a weak Kerr nonlinear with strength $K=3$ kHz, we apply a two-phonon drive to the mechanical mode satisfying $\Omega_{p}/\delta_{a}=\tanh(2r)$ with $r=1.5$. In this case, the anharmonicity and decoherence rate are then enhanced to $\sim 1.8$ MHz and $\sim 13.5$ kHz, where we have $\gamma\ll\alpha$, such that the mechanical resonator can be seen as a two-level system with a frequency $\omega_{b}=4.2$ MHz. Mechanical resonators with weak nonlinearity have been reported in nanomechanical systems such as cantilever, membranes or carbon nanotubes [78, 79, 80, 81, 82, 83]. In addition, our proposed scheme is also applicable to the systems whose nonlinearities are introduced by coupling to two energy level systems such as superconducting qubits [84, 85, 40, 52].
We next consider using this mechanical squeezed Fock qubit to sense weak force filed $kx^{2}/2$. For a typical example of carbon nanotube resonators with frequency $\omega_{a}$, the mass is $m=10^{-21}$ kg, we have the zero-point fluctuation $x_{0}=4\times 10^{-12}$ m. Under a decoherence rate of $\gamma_{0}/2\pi=3$ kHz, the calculated sensitivity reaches $\delta k_{\text{min}}=4.71\times 10^{-10}$ $\text{N}\cdot\text{m}^{-1}/\text{Hz}$. This represents an order-of-magnitude improvement over the undriven case with identical parameters ($k_{0}=2.11\times 10^{-9}$ $\text{N}\cdot\text{m}^{-1}/\text{Hz}$), and the enhancement becomes more pronounced at larger values of $r$.
Conclusion.—In conclusion, we have shown that, bt applying a two-phonon drive to mechanical resonators with weak Kerr nonlinearities, one can obtain an exponentially enhanced, tunable anharmonicities. Though the decoherence is also expeonentialy enhanced, the main benefit is that the enhancement of the former is much greater than that of the latter. This scheme enables us to encode a mechanical qubit in the first two squeezed-Fock phonon states. The mechanical qubit can couple to a variety of forces and thus be used to sense weak signals. Compared to the traditional mechanical Fock qubit, the sensitivity using the squeezed-Fock qubit has been exponentially promoted. The proposed squeezed-Fock state encoding is general and can be applied to other bosonic systems like photons and magnons.
Acknowledgements.
P.B.L. is supported by the National Natural Science Foundation of China under Grants No. W2411002 and No. 12375018. J.H.A. is supported by the National Natural Science Foundation of China (Grants No. 12275109 and No. 12247101) and the Innovation Program for Quantum Science and Technology of China (Grant No. 2023ZD0300904).